In how many ways can a committee of three be chosen from 4 married couples if


a) All are equally eligible?


b) One particular man must be on the committee?

1
Expert's answer
2022-04-19T01:57:12-0400

a) "N=C_8^3=\frac{8!}{3!5!}=56."

b) "N=C_7^2=\frac{7!}{2!5!}=21."